How would I find the x-intercepts for the quadratic equations:
6x^2 - 7x + 2 = 0
5x^2 – 21x + 4 = 0
4x^2 – 3x - 1 = 0

Answer by stanbon(75887)

How would I find the x-intercepts for the quadratic equations:
Factor and solve:
=========================
6x^2 - 7x + 2 = 0
(6x^2 -3x-4x +2 = 0
3x(2x-1)-2(2x-1) = 0
(2x-1)(3x-2) = 0
x = 1/2 or x = 2/3
==============================

5x^2 – 21x + 4 = 0
5x^2-20x-x+4 = 0
5x(x-4)-(x-4) = 0
(x-4)(5x-1) = 0
x = 4 or x = 1/5
==============================
4x^2 – 3x - 1 = 0
4x^2 -4x+x - 1 = 0
4x(x-1)+(x-1) = 0
(x-1)(4x+1) = 0
x = 1 or x = -1/4
==============================
Cheers,
Stan H.